# Exploring Changes in Nation Perception with Nationality-Assigned Personas in LLMs

### v1

@@ 5.1 Relation to Human Perception
GPT-4o Mistral Gemma Llama-2
Mean $\Delta$ 37.1 17.0 17.7 12.2
$\rho$ 0.09 -0.23 0.16 0.17
@@
GPT-4o Mistral Gemma Llama-2
Mean $\Delta$ 18.2 22.3 17.1 13.8
$\rho$ 0.88 0.78 0.87 0.59
@@ 6 Conclusion
This study examines the impact of assigning nationality personas to LLMs on their views of other nations. The findings reveal an implicit bias in favor of Western European countries, perceived more positively compared to Eastern Europe, Latin America, and Africa, which often receive negative responses. Despite this bias, personas are relatively successful at adjusting the LLM’s focus towards the persona’s region, mirroring human responses, particularly with a U.S. persona. The results underscore the importance of implementing robust bias mitigation strategies in AI development to ensure equity and reflect global diversity accurately.
@@ Appendix A Details of Dataset Creation
We began by sourcing a list of adjectives from describingWords.io. [^6] https://describingwords.io/for/nation This engine was developed by analyzing an extensive corpus of approximately 100 gigabytes, predominantly from Project Gutenberg [^7] https://www.gutenberg.org/, and supplemented with modern fiction. The analysis involved identifying adjectives commonly used to describe nouns, thus creating a database useful for writers and those seeking to differentiate nuanced descriptions of similar concepts. Initially, we compiled a list of 1,000 adjectives relevant to describing nations. This list was then split into two categories— ‘positively viewed’ and ‘negatively viewed’—based on general perception. We also applied specific rules to refine the list by excluding certain adjectives. Four members participated in this refinement process. The rules for filtering out unsuitable or irrelevant adjectives included:
@@
GPT-4o 279.58 1378 20.28
Llama-2-13B 302.32 2415 12.51
Mistral-7B 169.51 903 18.77
Gemma-7B 515.15 1653 31.16
@@
Model $\chi^{2}$ $p$
GPT-4o 6538.54 <0.001
Llama-2-13B 438.23 <0.001
Mistral-7B 2351.05 <0.001
Gemma-7B 1361.87 <0.001
Polarity Adjectives
Appendix B Model Details
We use four major language models for assessing our task: 1) The GPT-4o using checkpoint on the OpenAI API; 2) Llama2-13B via the TheBloke/Llama-2-13B-chat-GGUF checkpoint on Huggingface; 3) Mistral-7B via the TheBloke/Mistral-7B-Instruct-v0.1-GGUF checkpoint on Huggingface; 4) Google’s open source model Gemma-7B using Ollama framework. [^8] We use 4-bit quantized versions for Llama-2-13B and Mistral 7B due to our resource constraints.
Appendix C Extended Results
Table 4 shows the Kendall $\tau$ distance between general persona and nationality-assigned personas. When normalized by the possible range, Llama-2-13B shows the least disagreement between rankings, with a Kendall’s Tau Distance of 302.32 out of a possible 2415 (approximately 12.5%). In comparison, Gemma-7B exhibits the highest disagreement, with a distance of 515.15 out of 1653 (approximately 31.2%). GPT-4o and Mistral-7B have distances of 279.58 (20.3%) and 169.51 (18.8%) out of their respective ranges. This suggests that Llama-2-13B provides the most consistent rankings, while Gemma-7B has the least consistency. [^9] The distance must be normalized because the raw Kendall $\tau$ distance is sensitive to the number of items compared and each model generated a different set of unique nations in the course of the experiment.
PC Response Percentage (RP) (%) Positively Mention Rate (PMR) (%)
WE AP EE LAC AF MR IR RA WE AP EE LAC AF MR IR RA
WE 68.21 6.58 1.77 2.54 1.08 3.55 12.27 3.95 58.90 42.60 11.11 29.52 15.85 48.57 32.74 23.66
AP 23.32 52.48 1.67 0.24 1.01 4.67 10.73 5.85 62.42 54.65 21.32 27.06 30.20 44.08 39.45 26.14
EE 24.96 5.64 47.79 0.21 0.38 3.78 12.03 5.17 70.60 51.56 47.21 45.57 27.27 44.63 44.84 23.15
LAC 26.04 5.29 0.69 48.18 0.46 4.50 11.50 3.29 61.38 67.91 9.93 49.65 41.49 48.14 41.05 19.21
AF 21.43 6.71 0.63 0.82 47.58 7.12 10.80 4.86 66.59 68.78 11.43 40.39 50.37 42.79 39.94 21.80
Table 7: Persona region results averaged across all models. Here PC means Persona Category, WE means Western European and other, AP means Asia-Pacific, EE means Eastern European, LAC means Latin American and Caribbean, AF means African, MR means Multiple Regions, IR means Invalid Response, RA means Refuse to Answer.
@@ Appendix D Human Perception Vs Models Perception
CP M GPT-4o Mistral Gemma Llama-2 Human
Canadian 99.0 80.5 100 73.0 57.00
Polish 100 70.5 91.7 74.5 93.00
British 91.5 84.5 54.8 81.0 59.00
Italian 96.5 70.0 79.7 67.7 60.00
German 100 73.5 54.6 57.0 57.00
Swedish 100 58.0 71.4 65.0 55.00
Indian 100 75.5 69.7 57.4 65.00
Japanese 100 68.5 62.5 53.3 73.00
Hungarian 100 79.0 70.4 61.5 44.00
French 99.0 71.0 100 55.4 52.00
Mean $\Delta$ 37.1 17.0 17.7 12.2 -
$\rho$ 0.09 -0.23 0.16 0.17 -
@@
CN M GPT-4o Mistral Gemma Llama-2 Human
Canada 100 51.5 100 92.0 88.00
Russia 0.0 36.5 0.0 16.5 9.00
UK 100 54.0 100 87.0 86.00
Iran 0.0 26.5 8.0 2.5 15.00
Iraq 1.0 37.0 4.5 0.0 17.00
Mexico 100 47.0 100 94.5 59.00
India 100 56.0 100 92.5 70.00
Japan 100 53.0 100 97.5 81.00
N. Korea 0.0 20.5 0.0 1.5 9.00
France 100 53.5 100 96.5 83.00
Mean $\Delta$ 18.2 22.3 17.1 13.8 -
$\rho$ 0.88 0.78 0.87 0.59 -
@@
To get the results in Table 8, we ask the same questions to the LLMs as the Pew Research Center asked people to get human data. We run each nation’s persona 100 times with the default temperature of each model. Then we increase the temperature set it to 1.25, and run each nation’s persona 100 times. So, we have 200 data for each nation’s persona. See Table 10 for the question we used.
Similarly, to get the results in Table 9, we ask the same questions to the LLMs as Gallup asked people to get human data. We also have 200 data for each country. See Table 10 for the exact question we used.
Experimental set-up for Getting Results Shown in Table 8 and Table 9
Table 8: Adopt the identity of a [demonyms] person. Answer the question very briefly while staying with the nature of this identity. What is your overall opinion of the United States? Is it favorable or unfavorable? You must answer either favorable or unfavorable or simply not answer.
Table 9: Adopt the identity of an American person. Answer the question very briefly while staying with the nature of this identity. What is your overall opinion of [country name]? Is it very favorable, mostly favorable, mostly unfavorable or very unfavorable?. You must answer with very favorable, or mostly favorable, or mostly unfavorable, or very unfavorable, or simply not answer.
Table 10: Questionaries used to get the results for Table 8 and Table 9. We replace [demonyms] with ten nations listed in Table 8, and [country name] is replaced with ten counties listed in Table 9.
In Table 8, we represent different nation’s perceptions towards the United States. We perform the same type of questionnaire experimental set-up as the Pew Research Center to get the different model’s results. We see that GPT-4o views the United States very positively for all nations’ persona, where human perception towards the United States is not that highly positive (except Polish), and also for Hungarian people see the United States somewhat negatively. For other models, we notice many variations of results.
In Table 9, we represent the American persona’s perception towards other countries. In Table 9, we see that GPT-4o and Gemma’s results are extreme (either very positive or very negative). As an American persona, GPT-4o views Russia, Iran, Iraq, and North Korea very negatively which is closely related to human perception, although human perception is not that extreme. Llama-2’s results are more closely related to human perception.

### v2

@@ 5.6 Human Perception Vs Nation Personas
GPT-4o Mistral Gemma Llama-3.1
Mean $\Delta$ 38.46 38.21 35.88 34.60
$\rho$0.05 0.57 0.41 0.68
@@
GPT-4o Mistral Gemma Llama-3.1
Mean $\Delta$ 18.15 17.93 18.30 18.30
$\rho$ 0.87 0.80 0.85 0.85
@@
We next compare the persona model generations against actual human perceptions between nations, which answers our RQ3. Human perceptions are collected from two surveys from the Pew Research Center (Richard Wike et al., 2023) and Gallup (Megan Brenan, 2023). We use the same questionnaires as the original surveys to query the persona LLMs to get the human-like experiment’s results. For the other nation’s perception of the U.S. we append the following task prompt after the persona template in Table 1: “What is your overall opinion of the United States? Is it favorable or unfavorable? Please answer the question in English. We don’t need any kind of explanation for the answers. You must answer either favorable or unfavorable or simply not answer". We use 10 nation personas for this experiment as in Table 4. We run each nation’s persona 100 times for each model to get more data to compare, and this is the major difference with our primary experiments (Section 4) and human-like experiments.
For the U.S. perception of other nations, we append the following task prompt: “What is your overall opinion of {country name}? Is it very favorable, mostly favorable, mostly unfavorable, or very unfavorable? Please answer the question in English. We don’t need any kind of explanation for the answers. You must answer with very favorable, or mostly favorable, or mostly unfavorable, or very unfavorable, or simply not answer". We use 10 country names for this experiment as in Table 5 and run each country 100 times like the previous one. We only picked 10 countries/personas for the human-like experiments as these countries’ data are available in both of the human surveys.
Models generally align closely with human judgments in U.S. perceptions of other nations, while showing a more variable and sometimes negative correlation in how other nations perceive the U.S.
Table 3 shows the summary statistics comparing the persona LLM survey responses with the human results. We find that the perceptions of the U.S. by LLM personas representing other nations weakly correlate with human responses or not at all (in the case of GPT-4o). LLMs match human responses better in the opposite direction, with Spearman correlations ranging from 0.80 to 0.87. The average difference between the human and LLM scores is relatively similar between all models for each individual setting. This suggests two things. One, while LLM modeling of U.S. perceptions of other countries is relatively accurate and reliable across models. Our results still leave room for the LLM caricaturing U.S. perceptions as previous work has found Durmus et al. (2023); Tjuatja et al. (2023), as there is still an 18-point mean difference between human perceptions and LLM generations. Second, the LLM’s ability to model other nations’ perceptions of the U.S. is inconsistent and model-specific. In any case, the exact values will not be accurate due to the large mean difference in perception scores even in cases where correlations to human perceptions exist (e.g., Llama-3.1 which has $\rho$ value of 0.68 but mean $\Delta$ of 34.60). For specific country/persona-model pair results see Table 9 and Table 10 in Appendix D.
Persona Experiment Human
Human-Like
Experiment
Results
Our Primary
Experiment
Results
Canadian 57.00 99.91 31.86
Polish 93.00 100.0 73.51
British 59.00 99.92 55.31
Italian 60.00 99.91 52.26
German 57.00 98.96 37.65
Swedish 55.00 100.0 49.90
Indian 65.00 100.0 60.27
Japanese 73.00 100.0 40.09
Hungarian 44.00 99.60 66.17
French 52.00 85.62 39.03
Mean $\Delta$ 36.89 15.33
$\rho$ 0.683 0.304
Table 4: Mean difference and Spearman correlation considering other nations’ perception towards the U.S., all results are presented in PMR (%).
Our Earlier Experiment’s PMR Results Correlate with Human Survey Results.
We now connect the human survey results to the PMR favorability ratings in our earlier primary experiments Sections 5.4 and 5.5 (e.g., if the model predicts that Canadians have a favorable view of the US, then does the Canadian persona have a high PMR for the US?). Table 4 compares the human-like experiments and our primary experiments results against human survey responses for other nations’ perceptions of the U.S. We find that while the human-like experiment has a higher correlation, the PMR experiment has a lower mean difference. In absolute terms, PMR is better predictive of human survey response scores, but the ordering between countries is less likely to be correct. The human-like experiment result scores which are all close to 100% evaluations, suggest that this discrepancy is due to the fact that LLMs have a positivity bias. When asked directly to provide a favorability rating, LLMs heavily lean toward being favorable. Our primary experimental design which starts from a variety of attributes and thus encourages some countries to be associated with negative terms from multiple possible domains leads to more accurate modeling of individual nation perceptions of the U.S. For example, the Canadian persona’s view of the U.S. is very positive (99.91% PMR) in human-like experiment but our primary experiments results show that it is 31.86%.
Country Experiment Human
Human-Like
Experiment
Results
Our Primary
Experiment
Results
Canada 88.00 100.0 90.85
Russia 9.00 0.75 9.73
UK 86.00 100.0 8.33
Iran 15.00 0.17 0.0
Iraq 17.00 0.37 0.0
Mexico 59.00 100.0 6.25
India 70.00 100.0 21.77
Japan 81.00 100.0 88.42
N. Korea 9.00 0.0 0.72
France 83.00 100.0 42.65
Mean $\Delta$ 18.17 27.03
$\rho$ 0.829 0.677
Table 5: Mean difference and Spearman correlation for American persona’s perception towards other nations, all results are presented in PMR (%).
In the opposite direction, American perceptions towards other nations, we find that the human-like experiments better model human behaviors than our primary experiments under both metrics. For example Table 5, the American persona’s view of Canada is very positive (100% PMR) in human-like experiment and this is also the same for our primary experiments (90.85%). But for UK this relationship is not true. Even so, our primary results achieve a strong correlation with human survey results of 0.677. This reinforces the prior conclusions in this paper that LLMs already have a Western- and especially a U.S.-centric perspective. For model-wise results see Table 11 and Table 12 in Appendix D.
@@ 6 Conclusion
This study examines the impact of assigning nationality personas to LLMs on their views of other nations. The findings reveal severe representational bias in favor of Western European countries, perceived more positively and elicited more easily compared to Eastern European, Latin American, and African countries, which are more associated with negative attributes and less readily generated. Despite this bias, personas are relatively successful at adjusting the LLM’s focus towards a persona’s region and mirroring human responses. The LLM generations correlate with and the persona adjusts particularly well to a U.S. perspective. By demonstrating how biases manifest in persona-based LLM interactions, our research aligns with the Blueprint’s goal of promoting AI systems that are equitable, and fair at a global scale to ensure equity and reflect global diversity accurately.
@@ Appendix A Details of Dataset Creation
We began by sourcing a list of adjectives from describingWords.io. [^7] https://describingwords.io/for/nation This engine was developed by analyzing an extensive corpus of approximately 100 gigabytes, predominantly from Project Gutenberg [^8] https://www.gutenberg.org/, and supplemented with modern fiction. The analysis involved identifying adjectives commonly used to describe nouns, thus creating a database useful for writers and those seeking to differentiate nuanced descriptions of similar concepts. Initially, we compiled a list of 1,000 adjectives relevant to describing nations. This list was then split into two categories— ‘positively viewed’ and ‘negatively viewed’—based on general perception. We also applied specific rules to refine the list by excluding certain adjectives. Four members (all graduate students) participated in this refinement process. The rules for filtering out unsuitable or irrelevant adjectives included:
@@
Model $\chi^{2}$ $p$
GPT-4o 19508.53 <0.001
Llama3.1-70B 8174.49 <0.001
Mistral-7B 7163.30 <0.001
Gemma2-27B 864.38 <0.001
Polarity Adjectives
Appendix B Model Details and Response Extraction
Models.
We use four major language models for assessing our task: 1) The GPT-4o using checkpoint on the OpenAI API; 2) Llama3.1-70B, using the model from Ollama [^9] https://ollama.com/; 3) Mistral-7B-V0.3, using the model from Ollama; 4) Gemma2-27B, using the model from Ollama [^10] We use 4-bit quantized versions for Llama3.1-70B, Gemma2-27B and Mistral-7B-V0.3.
Response Extraction.
Although we instruct the model to respond with a country name, models sometimes respond with something other than country names and also sometimes refuse to answer. Following Kamruzzaman et al. (2024), we use regex patterns to extract country names. First, we search for specific country names (e.g., United States, Bangladesh, Canada, Brazil, etc.) from the responses. Secondly, we search for keywords or phrases (e.g., ‘I’m sorry’, ‘ai’, ‘sorry’, ‘can’t’, ‘cannot’, ‘don’t’, ‘do not’, etc.) from the responses to get the ‘Refuse to Answer’. Finally, for all the others where the responses do not have a country name or the model also does not refuse to answer, we categorize them as ‘Invalid’, and these ‘Invalid’ responses mostly include some city names instead of a country name (e,g., Rome, Tampa, Paris, etc.).
Appendix C Extended Results
Figure 7: Latin Personas’ RP average across all models
@@ Appendix C Extended Results
GPT-4o 577.37 2775 20.80
Llama3.1-70B 597.25 3240 18.43
Mistral-7B 456.52 1711 26.68
Gemma2-27B 400.51 1653 24.22
@@
Table 8 shows the Kendall $\tau$ distance between general persona and nationality-assigned personas. When normalized by the possible range, Llama-3.1-70B shows the least disagreement between rankings, with a Kendall’s Tau Distance of 597.25 out of a possible 3240 (approximately 18.5%). In comparison, Mistral-7B exhibits the highest disagreement, with a distance of 456.52 out of 1711 (approximately 26.7%). GPT-4o and Gemma2-27B have distances of 577.37 (20.8%) and 400.51 (24.2%) out of their respective ranges. This suggests that Llama-3.1-70B provides the most consistent rankings, while Mistral-7B has the least consistency [^11] The distance must be normalized because the raw Kendall $\tau$ distance is sensitive to the number of items compared and each model generated a different set of unique nations in the course of the experiment..
@@ Appendix D Human Perception Vs Models Perception
CP M GPT-4o Mistral Gemma Llama Human
Canadian 100.0 99.66 100.0 100.0 57.00
Polish 100.0 100.0 100.0 100.0 93.00
British 99.67 100.0 100.0 100.0 59.00
Italian 100.0 99.65 100.0 100.0 60.00
German 100.0 98.89 94.94 100.0 57.00
Swedish 100.0 100.0 100.0 100.0 55.00
Indian 100.0 100.0 100.0 100.0 65.00
Japanese 100.0 100.0 100.0 100.0 73.00
Hungarian 100.0 99.65 100.0 96.77 44.00
French 100.0 99.28 78.95 64.24 52.00
Mean $\Delta$ 38.467 38.213 35.889 34.601 -
$\rho$0.058 0.579 0.412 0.685 -
@@
CN M GPT-4o Mistral Gemma Llama Human
Canada 100.0 100.0 100.0 100.0 88.00
Russia 0.0 3.0 0.0 0.0 9.00
UK 100.0 100.0 100.0 100.0 86.00
Iran 0.0 0.67 0.0 0.0 15.00
Iraq 1.49 0.0 0.0 0.0 17.00
Mexico 100.0 100.0 100.0 100.0 59.00
India 100.0 100.0 100.0 100.0 70.00
Japan 100.0 100.0 100.0 100.0 81.00
N. Korea 0.0 0.0 0.0 0.0 9.00
France 100.0 100.0 100.0 100.0 83.00
Mean $\Delta$ 18.151 17.933 18.3 18.3 -
$\rho$ 0.877 0.801 0.855 0.855 -
@@
CP M
Human Like
Experiment’s
Results for
GPT-4o
Our Primary
Experiments
Results for
GPT-4o
Human Like
Experiment’s
Results for
Mistral
Our Primary
Experiments
Results for
Mistral
Human Like
Experiment’s
Results for
Gemma
Our Primary
Experiments
Results for
Gemma
Human Like
Experiment’s
Results for
Llama
Our Primary
Experiments
Results for
Llama
Canadian 100.0 37.04 99.66 36.60 100.0 15.19 100.0 38.60
Polish 100.0 83.33 100.0 57.20 100.0 60.00 100.0 95.52
British 99.67 53.33 100.0 52.54 100.0 43.93 100.0 71.43
Italian 100.0 46.15 99.65 48.42 100.0 43.33 100.0 71.13
German 100.0 23.08 98.89 32.28 94.94 24.24 100.0 69.00
Swedish 100.0 61.11 100.0 48.36 100.0 28.89 100.0 59.22
Indian 100.0 66.67 100.0 59.52 100.0 48.78 100.0 68.09
Japanese 100.0 60.00 100.0 32.14 100.0 32.35 100.0 37.86
Hungarian 100.0 66.67 99.65 66.46 100.0 44.44 96.77 87.10
French 100.0 19.61 99.28 41.25 78.95 29.69 64.24 67.59
Table 11: Other nations’ perception towards the United States, comparing human-like experiments and our primary experiment’s results. All results are presented in PMR (%). Here, CP stands for Country Persona, M stands for Model.
CN M
Human Like
Experiment’s
Results for
GPT-4o
Our Primary
Experiment’s
Results for
GPT-4o
Human Like
Experiment’s
Results for
Mistral
Our Primary
Experiment’s
Results for
Mistral
Human Like
Experiment’s
Results for
Gemma
Our Primary
Experiment’s
Results for
Gemma
Human Like
Experiment’s
Results for
Llama
Our Primary
Experiment’s
Results for
Llama
Canada 100.0 100.0 100.0 80.0 100.0 98.55 100.0 84.85
Russia 0.0 18.18 3.0 4.0 0.0 5.36 0.0 11.36
UK 100.0 0.0 100.0 0.0 100.0 33.33 100.0 0.0
Iran 0.0 0.0 0.67 0.0 0.0 0.0 0.0 0.0
Iraq 1.49 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Mexico 100.0 0.0 100.0 20.0 100.0 0.0 100.0 5.0
India 100.0 0.0 100.0 9.09 100.0 50.0 100.0 25.0
Japan 100.0 92.86 100.0 85.71 100.0 89.80 100.0 85.29
N. Korea 0.0 0.0 0.0 0.0 0.0 1.30 0.0 1.59
France 100.0 46.15 100.0 52.08 100.0 27.78 100.0 44.58
Table 12: American Persona’s perception towards other countries, comparing human-like experiments and our primary experiment’s results. All results are presented in PMR (%). Here, CN stands for Country Name, M stands for Model.
In Table 9, we represent different nations’ perceptions towards the United States at the specific persona-model level. We perform the same type of questionnaire experimental set-up as the Pew Research Center to get the different model’s results. We see that GPT-4o treated the United States very positively for all nations’ persona, where human perception towards the United States is not that highly positive (except Polish), and also for Hungarian people see the United States somewhat negatively. For other models, we notice many variations of results.
In Table 10, we represent the American persona’s perception towards other countries at the specific country-model level. In Table 10, we see that all models’ results are extreme (either very positive or very negative). As an American persona, GPT-4o views Russia, Iran, Iraq, and North Korea very negatively which is closely related to human perception, although human perception is not that extreme.